# Online conformal inference for multi-step time series forecasting

### v1

@@ 3.2 Online multi-step weighted conformal prediction
Barber et al. (2023) propose a nonexchangeable conformal prediction procedure (NexCP) that generalizes the SCP method to allow for some sources of nonexchangeability. The core idea is that a higher weight should be assigned to a data point that is believed to originate from the same distribution as the test data. Note that NexCP assumes the weights are fixed and data-independent. When the data are exchangeable, NexCP offers the same coverage guarantees as SCP, while the coverage gap is characterized by the total variation between the swapped nonconformity score vectors when exchangeability is violated. Thus the coverage gap may be quite large in time series contexts.
@@
We choose fixed weights $w_{i}=b^{t+1-i}$, $b\in(0,1)$ and $i=t-t_{c}+1,\ldots,t$, for nonconformity scores on the corresponding calibration set. In this setting, weights decay exponentially as the nonconformity scores get older, akin to the rationale behind the simple exponential smoothing method in time series forecasting (Hyndman & Athanasopoulos 2021). Then for each $h\in[H]$, MWCP updates the $h$-step-ahead prediction interval:
@@
$\tilde{w}_{i}=\frac{w_{i}}{\sum_{i=t-t_{c}+1}^{t}w_{i}+1},\text{ for }i\in\{t-t_{c}+1,\ldots,t\}\quad\text{and}\quad\tilde{w}_{t+1}=\frac{1}{\sum_{i=t-t_{c}+1}^{t}w_{i}+1}.$
@@ 3.3 Multi-step adaptive conformal prediction
Next we extend the adaptive conformal prediction (ACP) method proposed by Gibbs & Candès (2021) to address multi-step time series forecasting, introducing the multi-step adaptive conformal prediction (MACP) method. Assuming that $\beta\mapsto\mathbb{P}(y_{t+h}\in\hat{\mathcal{C}}_{t+h|t}^{\beta})$ is continuous and non-increasing, with $\mathbb{P}(y_{t+h}\in\hat{\mathcal{C}}_{t+h|t}^{0})=1$ and $\mathbb{P}(y_{t+h}\in\hat{\mathcal{C}}_{t+h|t}^{1})=0$, an optimal value $\alpha_{t+h|t}^{*}\in[0,1]$ exists such that the realised miscoverage rate of the corresponding prediction interval closely approximates the nominal miscoverage rate $\alpha$. Specifically, for each $h\in[H]$, we iteratively estimate $\alpha_{t+h|t}^{*}$ by updating a parameter $\alpha_{t+h|t}$ through a sequential adjustment process
@@
Then the $h$-step-ahead prediction interval is computed using Equation (1) by setting $\alpha=\alpha_{t+h|t}$. Here, $\gamma>0$ denotes a fixed step size parameter, $\alpha_{2h|h}$ denotes the initial estimate typically set to $\alpha$, and $\mathrm{err}_{t|t-h}$ denotes the miscoverage event $\mathrm{err}_{t|t-h}=\mathbb{1}\left\{y_{t}\notin\hat{\mathcal{C}}_{t|t-h}^{\alpha_{t|t-h}}\right\}$.
Equation (2) indicates that the correction to the estimation of $\alpha_{t+h|t}^{*}$ at time $t+h$ is determined by the historical miscoverage frequency up to time $t$. At each iteration, we raise the estimate of $\alpha_{t+h|t}^{*}$ used for quantile estimation at time $t+h$ if $\hat{\mathcal{C}}_{t|t-h}^{\alpha_{t|t-h}}$ covered $y_{t}$, whereas we lower the estimate if $\hat{\mathcal{C}}_{t|t-h}^{\alpha_{t|t-h}}$ did not cover $y_{t}$. Thus the miscoverage event has a delayed impact on the estimation of $\alpha_{t+h|t}^{*}$ over $h$ periods, indicating that the correction of the $\alpha_{t+h|t}^{*}$ estimate becomes less prompt with increasing values of $h$. In particular, Equation (2) reduces to the update for ACP for $h=1$.
@@
Selecting the parameter $\gamma$ is pivotal yet challenging. Gibbs & Candès (2021) suggest setting $\gamma$ in proportion to the degree of variation of the unknown $\alpha_{t}^{*}$ over time. Several strategies have been proposed to avoid the necessity of selecting $\gamma$. For example, Zaffran et al. (2022) use an adaptive aggregation of multiple ACPs with a set of candidate values for $\gamma$ , determining weights based on their historical performance. Bastani et al. (2022) propose a multivalid prediction algorithm in which the prediction set is established by selecting a threshold from a sequence of candidate thresholds. However, both previous methods fail to promptly adapt to the local changes. To address this limitation, Gibbs & Candès (2024) suggest adaptively tuning the step size parameter $\gamma$ in an online setting, choosing an “optimal” value for $\gamma$ from a candidate set of values by assessing their historical performance.
The theoretical coverage properties of ACP suggest that a larger value for $\gamma$ generally results in less deviation from the target coverage. As there is no restriction on $\alpha_{t+h|t}$ and it can drift below 00 or above $1$, a larger $\gamma$ may lead to frequent output of null or infinite prediction sets in order to quickly adapt to the current miscoverage status.
@@ 3.4 Multi-step conformal PID control
We introduce multi-step conformal PID control method (hereafter MPID), which extends the PID method (Angelopoulos et al. 2023), originally developed for one-step-ahead forecasting. For each individual forecast horizon $h\in[H]$, the iteration of the $h$-step-ahead quantile estimate is given by
@@
where $\eta>0$ is a constant learning rate, and $r_{t}$ is a saturation function that adheres to the following conditions
@@
for constant $b,c>0$, and an admissible function $g$ that is sublinear, nonnegative, and nondecreasing. With this updating equation, we can obtain all required $h$-step-ahead prediction intervals using information available at time $t$. When $h=1$, Equation (3) simplifies to the PID update, which guarantees long-run coverage. More importantly, Equation (3) represents a specific instance of Equation (10) that we will introduce later, thereby ensuring long-run coverage for each individual forecast horizon $h$ according to Proposition 5.
The P control in MPID shows a delayed correction of the quantile estimate for a length of $h$ periods. The underlying intuition is similar to that of MACP: it increases (or decreases) the $h$-step-ahead quantile estimate if the prediction set at time $t$ miscovered (or covered) the corresponding realisation. MACP can be considered as a special case of the P control, while the P control has the ability to prevent the generation of null or infinite prediction sets after a sequence of miscoverage events.
The I control accounts for the cumulative historical coverage errors associated with $h$-step-ahead prediction intervals during the update process, thereby enhancing the stability of the interval coverage.
The D control involves $\hat{s}_{t+h|t}$ as the $h$-step-ahead forecast of the nonconformity score (i.e., the forecast error), produced by any suitable forecasting model (or “scorecaster”) trained using the $h$-step-ahead nonconformity scores available up to and including time $t$. This module provides additional benefits only when the scorecaster is well-designed and there are predictable patterns in the nonconformity scores; otherwise, it may increase variability in the coverage and prediction intervals.
@@ 4 Autocorrelated multi-step conformal prediction
On the other hand, in any non-trivial context, multi-step forecasts are correlated with each other. Specifically, the $h$-step-ahead forecast errors $e_{t+h|t}$ will be correlated with the forecast errors from the past $h-1$ steps, as forecast errors accumulate over the forecast horizon. However, no conformal prediction methods have taken this potential dependence into account in their methodological construction.
In this section, we will explore the properties of multi-step forecast errors and propose a novel conformal prediction method that accounts for their autocorrelations.
@@ 4.1 Properties of multi-step forecast errors
We assume that a time series $\{y_{t}\}_{t\geq 1}$ is generated by a general non-stationary autoregressive process given by:
@@
where $f_{t}$ is a nonlinear function in $d$ lagged values of $y_{t}$, the current value of the exogenous predictors, along with the preceding $k$ values, and $\varepsilon_{t}$ is white noise innovation with mean zero and constant variance. The sequence of model coefficients that parameterizes the function $f$ is restricted to ensure that the stochastic process is locally stationary.
@@
${e_{t+h|t}=\omega_{t+h}+\phi_{1}e_{t+h-1|t}+\cdots+\phi_{p}e_{t+h-p|t},}$ (6)
[…]
Proposition 1 suggests that the optimal $h$-step ahead forecast error, $e_{t+h|t}$, is serially correlated with the forecast errors from at most the past $h-1$ steps, i.e., $e_{t+1|t},\ldots,e_{t+h-1|t}$. However, the autocorrelation among errors associated with optimal forecasts can not be used to improve forecasting performance, as it does not incorporate any new information available when the forecast was made. If we could forecast the forecast error, then we could improve the forecast, indicating that the initial forecast couldn’t have been optimal.
The proof of Proposition 1, provided in Section A.1, suggests that, if $f_{t}$ is a linear autoregressive model, then the coefficients in Equation (6) are the linear coefficients of the optimal forecasting model. However, when $f_{t}$ takes on a more complex nonlinear structure, the coefficients become complicated functions of observed data and unobserved model coefficients.
[…]
Proposition 2 (MA$(h-1)$ process for $h$-step-ahead optimal forecast errors).
Let $\{y_{t}\}_{t\geq 1}$ be a time series generated by a general non-stationary autoregressive process as given in Equation (5), and assume that any exogenous predictors are known into the future. Then the forecast errors of optimal $h$-step-ahead forecasts follow an approximate MA($h-1$) process
${e_{t+h|t}=\omega_{t+h}+\theta_{1}\omega_{t+h-1}+\cdots+\theta_{h-1}\omega_{t+1}.}$ (7)
@@ 4.2 The AcMCP method
The AcMCP method updates the quantile estimate $q_{t}$ in an online setting to achieve the goal of long-run coverage. Specifically, the iteration of the $h$-step-ahead quantile estimate is given by
@@
for $h\in[H]$. Obviously, the AcMCP method can be viewed as a further extension of the PID method. Nevertheless, AcMCP diverges from PID with several innovations and differences.
First, we are no longer confined to predicting just one step ahead. Instead, we can make multi-step forecasts with accompanying theoretical coverage guarantees, constructing distribution-free prediction intervals for steps $t+1,\ldots,t+H$ based on available information up to time $t$.
[…]
@@ 4.3 Coverage guarantees
Let $\{s_{t+h|t}\}_{t\in\mathbb{N}}$ be any sequence of numbers in $[-b,b]$ for any $h\in[H]$, where $b>0$, and may be infinite. Assume that $r_{t}$ is a saturation function obeying Equation (4), for an admissible function $g$. Then the iteration $q_{t+h|t}=r_{t}\left(\sum_{i=h+1}^{t}(\mathrm{err}_{i|i-h}-\alpha)\right)$ satisfies
@@
Proposition 3 indicates that, for tmuch-less-$t\ll\infty$, increasing the forecast horizon $h$ tends to amplify deviations from the target coverage because $g(T-h)/(t-h)$ is non-increasing, given that the admissible function $g$ is sublinear, nonnegative, and nondecreasing. This is consistent with expectations, as extending the forecast horizon generally increases forecast uncertainty. As predictions extend further into the future, more factors contribute to variability and uncertainty. In this case, conformal prediction intervals may not scale perfectly with the increasing uncertainty, leading to a larger discrepancy between the desired and actual coverage.
@@
Proposition 4.
Let $\{s_{t+h|t}\}_{t\in\mathbb{N}}$ be any sequence of numbers in $[-b,b]$ for any $h\in[H]$, where $b>0$, and may be infinite. Then the iteration $q_{t+h|t}=q_{t+h-1|t-1}+\eta(\mathrm{err}_{t|t-h}-\alpha)$ satisfies
@@
More importantly, Proposition 3 is also adequate for establishing the coverage guarantee of the proposed AcMCP method given byEquation (8). Following Angelopoulos et al. (2023), we first reformulate Equation (8) as
@@
where $\hat{q}_{t+h|t}$ can be any function of the past observations $\{(\bm{x}_{i},y_{i})\}_{1\leq i\leq t}$ and quantile estimates $q_{i+h|i}$ for $i\leq t-1$. Taking $\hat{q}_{t+h|t}=q_{t+h-1|t-1}+\eta(\mathrm{err}_{t|t-h}-\alpha)+\tilde{e}_{t+h|t}$ will recover Equation (8). We can consider $\hat{q}_{t+h|t}$ as the forecast of the quantile $q_{t+h|t}$ based on available historical data. We then present the coverage guarantee for AcMCP given by Equation (10).
Proposition 5.
Let $\{\hat{q}_{t+h|t}\}_{t\in\mathbb{N}}$ be any sequence of numbers in $[-\frac{b}{2},\frac{b}{2}]$, and $\{s_{t+h|t}\}_{t\in\mathbb{N}}$ be any sequence of numbers in $[-\frac{b}{2},\frac{b}{2}]$, for any $h\in[H]$, $b>0$ and may be infinite. Assume that $r_{t}$ is a saturation function obeying Equation (4), for an admissible function $g$. Then the prediction intervals obtained by the AcMCP iteration given by Equation (10) yield the correct long-run coverage; i.e., $\lim_{T\rightarrow\infty}\frac{1}{T-h}\sum_{t=h+1}^{T}\mathrm{err}_{t|t-h}=\alpha$.

### v2

@@ 3.2 Online multi-step weighted conformal prediction
Barber et al. (2023) propose nonexchangeable conformal prediction procedure (NexCP) that generalizes SCP to certain nonexchangeable settings by assigning higher weights to observations believed to share the same distribution as the test data. NexCP assumes these weights are fixed and data-independent. Under exchangeability, NexCP retains the same coverage guarantees as SCP; when exchangeability is violated, the coverage gap is governed by the total variation distance between swapped nonconformity score vectors, which can be substantial in time series settings.
@@
For the subsequent empirical study, we choose fixed weights $w_{i}=b^{t+1-i}$, $b\in(0,1)$ and $i=t-t_{c}+1,\dots,t$, for nonconformity scores on the corresponding calibration set. In this setting, weights decay exponentially as the nonconformity scores get older, akin to the rationale behind the simple exponential smoothing method in time series forecasting (Hyndman & Athanasopoulos 2021). Then for each $h\in[H]$, MWCP updates the $h$-step-ahead prediction interval:
@@
$\tilde{w}_{i}=\frac{w_{i}}{\sum_{i=t-t_{c}+1}^{t}w_{i}+1},\text{ for }i\in\{t-t_{c}+1,\dots,t\}\quad\text{and}\quad\tilde{w}_{t+1}=\frac{1}{\sum_{i=t-t_{c}+1}^{t}w_{i}+1}.$
@@ 3.3 Multi-step adaptive conformal prediction
Next we extend the adaptive conformal inference (ACI, denoted as ACP hereafter) method proposed by Gibbs & Candès (2021) to address multi-step time series forecasting, introducing the multi-step adaptive conformal prediction (MACP) method. Assuming that $\beta\mapsto\mathbb{P}(y_{t+h}\in\hat{\mathcal{C}}_{t+h|t}^{\beta})$ is continuous and non-increasing, with $\mathbb{P}(y_{t+h}\in\hat{\mathcal{C}}_{t+h|t}^{0})=1$ and $\mathbb{P}(y_{t+h}\in\hat{\mathcal{C}}_{t+h|t}^{1})=0$, an optimal value $\alpha_{t+h|t}^{*}\in[0,1]$ exists such that the realised miscoverage rate of the corresponding prediction interval closely approximates the nominal miscoverage rate $\alpha$. Specifically, for each $h\in[H]$, we iteratively estimate $\alpha_{t+h|t}^{*}$ by updating a parameter $\alpha_{t+h|t}$ through a sequential adjustment process
@@
Then the $h$-step-ahead prediction interval is computed using Equation (1) by setting $\alpha=\alpha_{t+h|t}$. Here, $\gamma>0$ denotes a fixed step size parameter, $\alpha_{2h|h}$ denotes the initial estimate typically set to $\alpha$, and $\mathrm{err}_{t|t-h}$ denotes the miscoverage event $\mathrm{err}_{t|t-h}=\mathbbm{1}\left\{y_{t}\notin\hat{\mathcal{C}}_{t|t-h}^{\alpha_{t|t-h}}\right\}$.
Equation (2) indicates that the correction to the estimation of $\alpha_{t+h|t}^{*}$ at time $t+h$ is determined by the historical miscoverage frequency up to time $t$. At each iteration, the estimate used for quantile calibration is increased if $\hat{\mathcal{C}}_{t|t-h}^{\alpha_{t|t-h}}$ covers $y_{t}$ and decreased otherwise. Therefore, the miscoverage event has a delayed impact on the estimation of $\alpha_{t+h|t}^{*}$ over $h$ periods, indicating that the correction of the $\alpha_{t+h|t}^{*}$ estimate becomes less prompt with increasing values of $h$. In particular, Equation (2) reduces to the update for ACP for $h=1$.
@@
Selecting the parameter $\gamma$ is pivotal yet challenging. Gibbs & Candès (2021) suggest setting $\gamma$ in proportion to the degree of variation of the unknown $\alpha_{t}^{*}$ over time. Several strategies have been proposed to avoid the necessity of selecting $\gamma$. For example, Zaffran et al. (2022) use an adaptive aggregation of multiple ACPs with a set of candidate values for $\gamma$, determining weights based on their historical performance. Bastani et al. (2022) propose a multivalid prediction algorithm in which the prediction set is established by selecting a threshold from a sequence of candidate thresholds. Both methods rely on update schemes that place substantial weight on older historical data, which may impede rapid adaptation to abrupt changes (Gibbs & Candès 2024). Therefore, Gibbs & Candès (2024) propose an alternative expert selection scheme that adaptively tunes the step-size parameter over time and places greater emphasis on more recent observations by construction, enabling faster responses to sudden environmental shifts.
The theoretical coverage properties of ACP suggest that a larger value for $\gamma$ generally results in less deviation from the target coverage. As there is no restriction on $\alpha_{t+h|t}$, and it can drift below 0 or above $1$, a larger $\gamma$ may lead to frequent output of null or infinite prediction sets in order to quickly adapt to the current miscoverage status.
@@ 3.4 Multi-step conformal PID control
We introduce multi-step conformal PID control method (hereafter MPID), which extends the PID method (Angelopoulos et al. 2023), originally developed for one-step-ahead forecasting. For each individual forecast horizon $h\in[H]$, the estimated $1-\alpha$ quantile of the $h$-step-ahead score at time $t$ is updated iteratively as
@@
where $\eta>0$ is a constant learning rate, and $r_{t}$ is a saturation function that adheres to
@@
for constant $b,c>0$, and an admissible function $g$ that is sublinear, nonnegative, and nondecreasing. With this updating equation, we can obtain all required $h$-step-ahead prediction intervals using information available at time $t$. When $h=1$, Equation (3) simplifies to the PID update, which guarantees long-run coverage. More importantly, Equation (3) represents a specific instance of Equation (10) that we will introduce later, thereby ensuring long-run coverage for each individual forecast horizon $h$ according to Corollary 2.
The “P” control shows an $h$-period delay in updating the quantile estimate. The underlying intuition is similar to that of MACP: it increases (or decreases) the $h$-step-ahead quantile estimate if the prediction set at time $t$ miscovered (or covered) the corresponding realization. MACP can be considered as a special case of the P control, while the P control has the ability to prevent the generation of null or infinite prediction sets after a sequence of miscoverage events.
The “I” control captures cumulative historical coverage errors associated with $h$-step-ahead prediction intervals during updates, thereby enhancing the stability of the interval coverage.
The “D” control involves $\hat{s}_{t+h|t}$ as the $h$-step-ahead forecast of the nonconformity score (i.e., the forecast error), produced by any suitable forecasting model (or “scorecaster”, see Section 4 for a detailed discussion) trained using the $h$-step-ahead nonconformity scores available up to and including time $t$. The effectiveness of the D control depends on the ability of the scorecaster to capture systematic and predictable patterns in the nonconformity scores, such as temporal dependence or conditional heteroskedasticity. Only when such structure is present and adequately modeled, this module provides additional benefits; otherwise, it may increase variability in the coverage and prediction intervals.
The MPID method will be improved in Section 4 by proposing the AcMCP method, which replaces the D-control component of MPID with a structured mechanism that explicitly captures autocorrelation in multi-step forecast errors. As a result, the long-run coverage guarantees established for AcMCP in Section 4.3 also apply to MPID.
@@ 4 Autocorrelated multi-step conformal prediction
On the other hand, in any non-trivial context, multi-step forecasts are correlated with each other. Specifically, the $h$-step-ahead forecast errors $e_{t+h|t}$ are correlated with the forecast errors from the previous h1h-1 steps, as errors propagate through the recursive forecasting procedure over the forecast horizon. However, no conformal prediction methods have taken this potential dependence into account in their methodological construction.
In this section, we will explore the properties of multi-step forecast errors and propose a novel conformal prediction method that accounts for their autocorrelations while providing theoretical long-run coverage guarantees.
@@ 4.1 Properties of multi-step forecast errors
We assume that a time series $\{y_{t}\}_{t\geq 1}$ follows a general non-stationary autoregressive process:
@@
where $f_{t}$ is a nonlinear function in $d$ lagged values of $y_{t}$, the current value of the exogenous predictors, along with the preceding $k$ values, and $\varepsilon_{t}$ is white noise innovation with mean zero and constant variance. The sequence of model coefficients that parameterizes the function $f$ is restricted to ensure that the stochastic process is locally stationary (Dahlhaus 2012).
@@
$e_{t+h|t}=\omega_{t+h}+\phi_{1}e_{t+h-1|t}+\dots+\phi_{p}e_{t+h-p|t},$} (6)
where $p=\min\{d,h-1\}$, and $\omega_{t}$ is white noise. The approximation is obtained in the proof via a first-order Taylor series expansion of the forecasting function, where higher-order remainder terms are neglected. Therefore, the optimal $h$-step-ahead forecast errors are at most serially correlated to lag (h1)(h-1).
Proof.
See Appendix A1. ∎
Proposition 1 suggests that the optimal $h$-step ahead forecast error, $e_{t+h|t}$, is serially correlated with the forecast errors from at most the past h1h-1 steps, i.e., $e_{t+1|t},\dots,e_{t+h-1|t}$. However, the autocorrelation among errors associated with optimal forecasts can not be used to improve forecasting performance, as it does not incorporate any new information available when the forecast was made. If we could forecast the forecast error, then we could improve the forecast, indicating that the initial forecast couldn’t have been optimal.
The proof of Proposition 1, provided in Appendix A1, suggests that, if $f_{t}$ is a linear autoregressive model, then the coefficients in Equation (6) are the linear coefficients of the optimal forecasting model. However, when $f_{t}$ takes on a more complex nonlinear structure, the coefficients become complicated functions of observed data and unobserved model coefficients.
It is well-established in the forecasting literature that, for a zero-mean covariance-stationary time series, the optimal linear least-squares forecasts have $h$-step-ahead errors that are at most MA(h1)(h-1) process (Harvey et al. 1997), a property that can be derived using Wold’s representation theorem. The statement can be extended to time series generated by a general non-stationary autoregressive process (Sommer 2023), as described in Proposition 2. We provide the proof of this proposition in Appendix A2 based on Proposition 1.
Proposition 2 (MA(h1)(h-1) process for $h$-step-ahead optimal forecast errors).
Let $\{y_{t}\}_{t\geq 1}$ be a time series generated by a general non-stationary autoregressive process as given in Equation (5), and assume that any exogenous predictors are known into the future. Under the approximation in Proposition 1, the forecast errors of optimal $h$-step-ahead forecasts follow an approximate MA(h1h-1) process
$e_{t+h|t}=\omega_{t+h}+\theta_{1}\omega_{t+h-1}+\dots+\theta_{h-1}\omega_{t+1}.$} (7)
@@ 4.2 The AcMCP method
The AcMCP method updates the quantile estimate $q_{t}$ in an online setting to achieve the goal of long-run coverage. Specifically, the iteration of the $h$-step-ahead quantile estimate for the score is
@@
for $h\in[H]$. Obviously, the AcMCP method can be viewed as a further extension of the PID method by Angelopoulos et al. (2023). Nevertheless, AcMCP diverges from PID with several innovations and differences.
First, we are no longer confined to predicting just one step ahead. Instead, we can make multi-step forecasts with accompanying theoretical coverage guarantees, constructing distribution-free prediction intervals for steps $t+1,\dots,t+H$ based on available information up to time $t$.
Additionally, in AcMCP, $\tilde{e}_{t+h|t}$ is a forecast combination of two simple models: one being an MA(h1)(h-1) model trained on the $h$-step-ahead forecast errors available up to and including time $t$ (i.e. $e_{1+h|1},\dots,e_{t|t-h}$), and the other a linear regression model trained by regressing $e_{t+h|t}$ on forecast errors from past steps (i.e. $e_{t+h-1|t},\dots,e_{t+1|t}$). Unlike MPID, which treats multi-step prediction problems independently across horizons, AcMCP performs multi-step conformal prediction recursively, explicitly accounting for the serial dependence among multi-step forecast errors. Importantly, the role of $\tilde{e}_{t+h|t}$ is not to forecast the nonconformity scores themselves, but to incorporate the autocorrelation structure of multi-step forecast errors into the construction of the resulting prediction intervals.
@@ 4.3 Coverage guarantees
Let $\{s_{t+h|t}\}_{t\in\mathbb{N}}$ be any sequence of numbers in [b,b][-b,b] for any $h\in[H]$, where $b>0$, and may be infinite. Assume that $r_{t}$ is a saturation function obeying Equation (4), for an admissible function $g$. Then the iteration $q_{t+h|t}=r_{t}\left(\sum_{i=h+1}^{t}(\mathrm{err}_{i|i-h}-\alpha)\right)$ satisfies
@@
Proof.
See Appendix A3. ∎
When $h=1$, Proposition 3 reduces to Proposition 2 of Angelopoulos et al. (2023). While Angelopoulos et al. (2023) only consider one-step-ahead forecasting, Proposition 3 extends their result to the multi-step setting and provides an explicit upper bound on the coverage gap. For finite $T$, Proposition 3 indicates that increasing the forecast horizon $h$ amplifies in-sample deviations from the target coverage because g(Th)/(Th)g(T-h)/(T-h) is non-increasing, given that the admissible function $g$ is sublinear, nonnegative, and nondecreasing. Importantly, this reflects finite-sample behavior rather than a breakdown of the asymptotic coverage guarantee. As expected, forecast uncertainty increases with the horizon, since predictions further into the future are affected by more sources of variability. In finite samples, conformal prediction intervals may not fully scale with this growing uncertainty, leading to a larger gap between nominal and empirical coverage.
@@
Corollary 1.
Let $\{s_{t+h|t}\}_{t\in\mathbb{N}}$ be any sequence of numbers in [b,b][-b,b] for any $h\in[H]$, where $b>0$, and may be infinite. Then the iteration $q_{t+h|t}=q_{t+h-1|t-1}+\eta(\mathrm{err}_{t|t-h}-\alpha)$ satisfies
@@
Proof.
See Appendix A4. ∎
When $h=1$, Corollary 1 reduces to Proposition 1 of Angelopoulos et al. (2023). More generally, Corollary 1 extends this result to the multi-step setting and establishes an explicit upper bound on the finite-sample coverage gap. The boundedness assumption on the nonconformity scores is imposed solely for theoretical analysis, ensuring control over the accumulation of miscoverage errors in the quantile tracking recursion. When the update is unraveled, the quantile tracker can be interpreted as an error integrator that accumulates past deviations of the empirical coverage from the target level. Bounded scores guarantee the stability of this accumulation and enable explicit bounds on the finite-sample coverage gap. Importantly, the bound itself is not required in practice, and the update proceeds agnostically and remains valid as long as the scores are finite.
More importantly, Proposition 3 is also adequate for establishing the coverage guarantee of AcMCP given by Equation (8). We first reformulate Equation (8) as
@@
where $\hat{q}_{t+h|t}$ can be any function of the past observations $\{(\bm{x}_{i},y_{i})\}_{1\leq i\leq t}$ and quantile estimates $q_{i+h|i}$ for $i\leq t-1$. Taking $\hat{q}_{t+h|t}=q_{t+h-1|t-1}+\eta(\mathrm{err}_{t|t-h}-\alpha)+\tilde{e}_{t+h|t}$ will recover Equation (8). We can consider $\hat{q}_{t+h|t}$ as the forecast of the quantile $q_{t+h|t}$ based on available historical data. We then present the long-run coverage guarantee for AcMCP given by Equation (10).
Corollary 2.
Let $\{\hat{q}_{t+h|t}\}_{t\in\mathbb{N}}$ be any sequence of numbers in $\frac{b}{2},\frac{b}{2}]$, and $\{s_{t+h|t}\}_{t\in\mathbb{N}}$ be any sequence of numbers in $\frac{b}{2},\frac{b}{2}]$, for any $h\in[H]$, $b>0$ and may be infinite. Assume that $r_{t}$ is a saturation function obeying Equation (4), for an admissible function $g$. Then the prediction intervals obtained by the AcMCP iteration given by Equation (10) yield the correct long-run coverage; i.e., $\lim_{T\rightarrow\infty}\frac{1}{T-h}\sum_{t=h+1}^{T}\mathrm{err}_{t|t-h}=\alpha$.
Proof.
See Appendix A5. ∎